# Matrix-Driven Instant Review: Confident Detection and Reconstruction of LLM Plagiarism on PC

### v2

@@ 2 Preliminaries
Our work primarily builds upon the foundation of matrix analysis. To avoid excessive technicality or verbosity, we only introduce the key tools and properties used in this paper, potentially omitting or abbreviating proofs. Interested readers are encouraged to refer to [21] for a comprehensive overview of matrix analysis.
2.1 Matrix Analysis: Singular Value and Polar Decompositions
@@
Polar Decomposition. The polar decomposition of $A$ has two common but distinct forms: the left decomposition $A=PW$ and the right decomposition $A=WQ$, where $P=(AA^{T})^{1/2}$ and $Q=(A^{T}A)^{1/2}$ are symmetric positive semidefinite matrices, and $W\in\mathbb{R}^{m\times n}$ (with orthonormal columns) is shared between both decompositions. We define the orthogonal part of $A$ as $\mathrm{Ortho}(A):=W$. When $A$ is full-rank, the orthogonal part of $A$ is unique.
@@
$W=UV^{T},$
where $U$ and $V$ are derived from the SVD of $A$. The symmetric factors satisfy $P=USU^{T}$ and $Q=VSV^{T}$. However, when $A$ is not invertible (with many singular values either exactly zero or close to zero), the computation of $W=UV^{T}$ becomes ill-conditioned.
SVD and Spectral Calculus. Assume $A\in\mathbb{R}^{m\times n}$ has SVD $A=USV^{T}$. Then:
$AA^{T}=U(SS^{T})U^{T},\quad A^{T}A=V(S^{T}S)V^{T}.$
@@
$f(AA^{T})A$
$=Uf(SS^{T})U^{T}USV^{T}=Uf(SS^{T})SV^{T}$
$=Ug(S)V^{T}$
$=USf(S^{T}S)V^{T}=USV^{T}Vf(S^{T}S)V^{T}$
$=Af(A^{T}A).$
In fact, $g(S)$ can represent any odd polynomial. For any odd function $G$, we may find a polynomial $f$ such that $f(x)$ agrees with $G(x)/x$ on all nonzero diagonal entries of $SS^{T}$. In particular, setting $G(x)=\mathrm{sign}(x)$, this spectral calculus yields the most “sensible” orthogonal part of $A$. We adopt the methods in [3] for a fast and relatively accurate implementation.
@@
$\max_{\{X|X^{T}X=I\}}\left(\operatorname{Tr}(AX^{T})\right).$
@@
  • Schatten $p$-Norm [37]: $\lVert A\rVert_{p}=\sqrt[p]{\sum_{i}\sigma_{i}^{p}(A)}$.
@@
To prove this, note that $\mathrm{RMSNorm}(y)$ is always a constant multiple ($\sqrt{n}$) of a unit vector when $y\neq 0$. Thus, $U$ maps all unit vectors to unit vectors. By linearity, this implies $\lVert xU\rVert=\lVert x\rVert$ and $xUU^{T}x^{T}=xx^{T}$. Taking $x$ over all eigenspaces of $UU^{T}$, we see that $1$ is the only eigenvalue of $UU^{T}$. Hence, $UU^{T}=\bm{1}_{n}$, and $U$ is orthogonal.
2.2 Large Deviation Theory
@@
Our research extensively involves random orthogonal matrices, particularly focusing on traces of such matrices. To obtain statistically meaningful $p$-values, traditional statistical $p$-tests become useless for our case, due to the interdependence of the elements within an orthogonal matrix. Instead, we rely on large deviation theory for deriving the $p$-value. For a comprehensive exposition of this theory, please refer to [11] and [4].
3 Methods
3.1 Problem Statement
Model plagiarism refers to the act of reusing part or all of another model’s weights without proper acknowledgment as required by the original model’s usage protocol. In this paper, we focus solely on weight reuse and do not address similarities arising from training data selection.
Given two large language models $A$ and $B$, with their parameters denoted as $\theta_{A}$ and $\theta_{B}$, we aim to determine whether $A$ and $B$ exhibit a relationship in their weights, based solely on the statistical properties of $\theta_{A}$ and $\theta_{B}$. The weight relationships we consider include, but are not limited to, the following cases:
  • Fine-tuning: Various kinds of SFT and RL included;
  • Upcycling [52, 19]: Continual pretraining with a larger model with weights initialized from a smaller base model;
  • Pruning [25, 1]: Removing certain channels or neurons of a base model to obtain a smaller model;
  • Comprehensive Obfuscation: A combination of all these cases above.
This task can be formulated as a binary classification problem $\Psi$, where the inputs are $(\theta_{A},\theta_{B})$ and the output is $\Psi(\theta_{A},\theta_{B})\in\{0,1\}$. Here, $1$ indicates that the two models are related, while 0 indicates they are unrelated.
Due to the vast number of parameters (on the order of billions) and the scarcity of samples, it is impractical to construct a machine learning framework that reliably determines whether two models share a weight relationship. Any learning-based approach to this task would likely suffer from extreme overfitting without producing trustworthy signals. To address this challenge, we leverage tools from statistics and random matrix theory.
Our algorithm focuses exclusively on detecting relationships between the weights $\theta_{A}$ and $\theta_{B}$. Consequently, all arguments presented below remain valid if $\theta_{A}$ and $\theta_{B}$ are swapped. For convenience, we may assume that model $A$ is the victim (reference model) and model $B$ is the suspect. However, it is important to note that our algorithm does not distinguish between the suspect and the victim. The identification of the suspect, victim, and plagiarism should be determined using additional evidence, particularly temporal and societal clues.
3.2 Architectural Assumptions
For simplicity, we assume that both model $A$ and model $B$ adopt a decoder-only Transformer architecture [47, 35], with word embeddings and unembeddings, Grouped Query Attention (GQA) [2], MLP layers with up and down projections [41], and RMSNorm layers [54].
Most other architectures can be analyzed within our framework, as GQA represents the most prevalent form of attention mechanism in modern Transformers. Both Multi-Head Attention (with an expansion rate of $1$) [47] and Multi-Query Attention (with one key-value head per layer) [40] can be treated as special cases of GQA.
While our method may also apply to other architectures, such as linear RNN variants, a specialized analysis tailored to each architecture is required for accurate results.
We consider the GQA architecture depicted in Figure 1, which is illustrated using Penrose tensor notation.
Figure 1: A Penrose notation of Grouped Query Attention (top) and an illustration of an equivalent model under certain transformations. Bias terms are omitted, and RMSNorms are not explicitly shown.
3.3 Equivalent Transformations
The following analysis has been partially derived in prior literature [53], but with significant gaps. For instance, previous work often omits critical components such as Rotary Positional Embeddings (RoPE), attention heads, RMSNorms, and other architectural details.
Assume that Model $A$ admits the weights depicted in the upper subfigure of Figure 1, and Model $B$ is equivalent to $A$ under certain transformations:
$\theta_{A}$ $=\{Q,K,V,O\}$
$\theta_{B}$ $=\{Q^{\prime},K^{\prime},V^{\prime},O^{\prime}\}.$
$Q^{\prime}$ $=U_{Q}QW_{Q},\quad K^{\prime}=U_{K}KW_{K},\quad V^{\prime}=U_{V}VW_{V},\quad O^{\prime}=W_{O}^{-1}OU_{O}^{-1},$
where $U_{Q},U_{K},U_{V},U_{O}$ and $W_{Q},W_{K},W_{V},W_{O}$ are transformation matrices applied to the original weights. These matrices represent modifications introduced during plagiarism or model adaptation.
this implies that $U_{Q}=U_{K}=U_{V}$, and they are all orthogonal matrices. In the context of Lie groups, we denote $U_{Q}=U_{K}=U_{V}\in\mathrm{O}(\mathrm{EmbDim})$. In contrast, $U_{O}$ can be less restrictive and may include a scaling factor: $U_{O}=\lambda U_{Q}$ ($\lambda\neq 0$). In Lie group notation, $U_{O}\in\mathrm{O}(\mathrm{EmbDim})\times\mathbb{R}$.
For the inner transformations $W_{Q},W_{K},W_{V},W_{O}$, the situation becomes more complex due to the presence of attention heads and nonlinear transformations (e.g., Softmax) across channels. Not all orthogonal transformations are permissible for inner transformations.
From the Penrose diagram in Figure 1, we propose a possible set of transformations for the inner matrices $W_{Q},W_{K},W_{V},W_{O}$. While we cannot prove that this set encompasses all possible transformations, it provides a plausible characterization:
$W_{Q}$ $=\mu\cdot P_{1}\otimes P_{2}\otimes S,$
$W_{K}$ $=\mu^{-1}\cdot P_{1}\otimes S,$
$V^{\prime}$ $=P_{1}\otimes H_{v},$
$W_{O}$ $=P_{1}\otimes P_{2}\otimes H_{v},$

### v3

@@ Appendix B Mathematical Background
Our work primarily builds upon the foundation of matrix analysis. To avoid excessive technicality or verbosity, we only introduce the key tools and properties used in this paper, potentially omitting or abbreviating proofs. Interested readers are encouraged to refer to horn2012matrix for a comprehensive overview of matrix analysis.
B.1 Matrix Analysis: Singular Value and Polar Decompositions
@@
Polar Decomposition. The polar decomposition of $A$ has two common but distinct forms: the left decomposition $A=PW$ and the right decomposition $A=WQ$, where $P=(AA^{\text{T}})^{1/2}$ and $Q=(A^{\text{T}}A)^{1/2}$ are symmetric positive semidefinite matrices, and $W\in\mathbb{R}^{m\times n}$ (with orthonormal columns) is shared between both decompositions. We define the orthogonal part of $A$ as $\mathrm{Ortho}(A):=W$. When $A$ is full-rank, the orthogonal part of $A$ is unique.
@@
$W=UV^{\text{T}},$
where $U$ and $V$ are derived from the SVD of $A$. The symmetric factors satisfy $P=USU^{\text{T}}$ and $Q=VSV^{\text{T}}$. However, when $A$ is not invertible (with many singular values either exactly zero or close to zero), the computation of $W=UV^{\text{T}}$ becomes ill-conditioned.
SVD and Spectral Calculus. Assume $A\in\mathbb{R}^{m\times n}$ has SVD $A=USV^{\text{T}}$. Then:
$AA^{\text{T}}=U(SS^{\text{T}})U^{\text{T}},\quad A^{\text{T}}A=V(S^{\text{T}}S)V^{\text{T}}.$
@@
$f(AA^{\text{T}})A$
$=Uf(SS^{\text{T}})U^{\text{T}}USV^{\text{T}}=Uf(SS^{\text{T}})SV^{\text{T}}$
$=Ug(S)V^{\text{T}}$
$=USf(S^{\text{T}}S)V^{\text{T}}=USV^{\text{T}}Vf(S^{\text{T}}S)V^{\text{T}}$
$=Af(A^{\text{T}}A).$
In fact, $g(S)$ can represent any odd polynomial. For any odd function $G$, we may find a polynomial $f$ such that $f(x)$ agrees with $G(x)/x$ on all nonzero diagonal entries of $SS^{\text{T}}$. In particular, setting $G(x)=\mathrm{sign}(x)$, this spectral calculus yields the most “sensible” orthogonal part of $A$. We adopt the methods from amsel2025polarexpressoptimalmatrix for a fast and relatively accurate implementation.
@@
$\max_{\{X|X^{\text{T}}X=I\}}\left(\operatorname{Tr}(AX^{\text{T}})\right).$
@@
  • Schatten $p$-Norm (schatten_cross_norms): $\lVert A\rVert_{p}=\sqrt[p]{\sum_{i}\sigma_{i}^{p}(A)}$.
@@
To prove this, note that $\mathrm{RMSNorm}(y)$ is always a constant multiple ($\sqrt{n}$) of a unit vector when $y\neq 0$. Thus, $U$ maps all unit vectors to unit vectors. By linearity, this implies $\lVert xU\rVert=\lVert x\rVert$ and $xUU^{\text{T}}x^{\text{T}}=xx^{\text{T}}$. Taking $x$ over all eigenspaces of $UU^{\text{T}}$, we see that $1$ is the only eigenvalue of $UU^{\text{T}}$. Hence, $UU^{\text{T}}=\bm{1}_{n}$, and $U$ is orthogonal.
B.2 Large Deviation Theory
@@
Our research extensively involves random orthogonal matrices, particularly focusing on traces of such matrices. To obtain statistically meaningful $p$-values, traditional statistical $p$-tests become useless for our case, due to the interdependence of the elements within an orthogonal matrix. Instead, we rely on large deviation theory for deriving the $p$-value. For a comprehensive exposition of this theory, please refer to DemboZeitouni1998 and anderson2010introduction.
Appendix C Preservation under Scaling and Permutation
Following Equation 1, we let $E$ and $E^{\prime}$ denote the vocabulary embeddings of model $A$ and $B$ respectively. We demonstrate that the final results remain invariant when an adversary applies both scaling and permutation to the matrix $E^{\prime}$. Specifically, we consider the transformed embedding $E^{\prime\prime}=\alpha E^{\prime}P^{\prime}$, where $P^{\prime}\in\mathrm{Perm}(\mathrm{n})$ is a permutation matrix, and $\alpha>0$ is a scaling factor.
Theorem 1.
Define the trace and the corresponding optimal permutation respectively as follows:
$T(E,E^{\prime})$ $=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left(E^{\prime\text{T}}E\right)^{\text{T}}\right),$
$\Pi(E,E^{\prime})$ $=\operatorname*{arg\,max}_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left(E^{\prime\text{T}}E\right)^{\text{T}}\right).$
Suppose an adversary applies a simultaneous scaling and permutation transformation to $E^{\prime}$, yielding $E^{\prime\prime}=\alpha E^{\prime}P^{\prime}$, where $P^{\prime}\in\mathrm{Perm}(\mathrm{n})$ is an arbitrary permutation matrix, and $\alpha>0$ is a positive scaling coefficient. In this case, the following properties hold:
$T(E,E^{\prime\prime})=T(E,E^{\prime}),\quad\Pi(E,E^{\prime\prime})=P^{\prime\text{T}}\Pi(E,E^{\prime}).$
Thus, such adversarial modifications are futile.
Proof.
We first analyze the trace $T(E,E^{\prime\prime})$:
$T(E,E^{\prime\prime})$ $=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left((E^{\prime\prime\text{T}}E)^{\text{T}}\right)\right)$
$=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left(((\alpha E^{\prime}P^{\prime})^{\text{T}}E)^{\text{T}}\right)\right)$
$=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left(\alpha E^{\text{T}}E^{\prime}P^{\prime}\right)\right).$
Since scaling by $\alpha>0$ does not affect the orthogonal factor in polar decomposition ($\mathrm{Ortho}$), we can simplify further:
$T(E,E^{\prime\prime})$ $=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}\left((E^{\text{T}}E^{\prime})P^{\prime}\right)\right)$
$=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P\cdot\mathrm{Ortho}(E^{\text{T}}E^{\prime})P^{\prime}\right)$
$=\max_{P\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(P^{\prime}P\cdot\mathrm{Ortho}\left(E^{\text{T}}E^{\prime}\right)\right).$
Let $R=P^{\prime}P$. Since $P^{\prime}$ is a fixed permutation matrix, as $P$ ranges over all permutations in $\mathrm{Perm}(\mathrm{n})$, so does $R$. Substituting $R$ into the expression, we obtain:
$T(E,E^{\prime\prime})$ $=\max_{R\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(R\cdot\mathrm{Ortho}\left(E^{\text{T}}E^{\prime}\right)\right)$
$=\max_{R\in\mathrm{Perm}(\mathrm{n})}\operatorname{Tr}\left(R\cdot\mathrm{Ortho}\left((E^{\prime\text{T}}E)^{\text{T}}\right)\right)$
$=T(E,E^{\prime}).$
Next, we examine the optimal permutation $\Pi(E,E^{\prime\prime})$. Recall that the trace achieves its maximum when $R=\Pi(E,E^{\prime})$. From the substitution $R=P^{\prime}P$, it follows that:
$P^{\prime}P=\Pi(E,E^{\prime})\!\implies\!P=P^{\prime-1}\Pi(E,E^{\prime})=P^{\prime\text{T}}\Pi(E,E^{\prime}).$
Thus, the optimal permutation for $E^{\prime\prime}$ is given by:
$\Pi(E,E^{\prime\prime})=P^{\prime\text{T}}\Pi(E,E^{\prime}).$
This completes the proof. ∎
Appendix D Uniform Distribution across Haar Measure
Theorem 2.
Let $E^{\prime}$ be an $m\times n$ random matrix initialized with i.i.d. Gaussian entries $\sim\mathcal{N}(0,\sigma^{2})$, and let $E$ be an $m\times n$ matrix such that $\mathrm{rank}(E)=n$ ($m\geq n$ because vocabulary size is always large). We know that $E^{\prime\text{T}}E$ is almost surely full-rank. Define the function $\mathrm{Ortho}(X)=X(X^{\text{T}}X)^{-1/2}$, which gives the orthogonal factor in the polar decomposition of $X$. Then $\mathrm{Ortho}(E^{\prime\text{T}}E)$ is uniformly distributed across $\mathrm{O}(n)$ with respect to the Haar measure.
Proof.
We proceed as follows:
Lemma. For any orthogonal matrix $U\in O(n)$ and any $n\times n$ matrix $X$,
$\mathrm{Ortho}(UX)=U\,\mathrm{Ortho}(X).$
Proof. By definition, $\mathrm{Ortho}(X)=X(X^{T}X)^{-1/2}$. Substituting $UX$ into the formula:
$\mathrm{Ortho}(UX)=(UX)\left((UX)^{\text{T}}(UX)\right)^{-1/2}=UX\left(X^{\text{T}}U^{\text{T}}UX\right)^{-1/2}=UX\left(X^{\text{T}}X\right)^{-1/2}=U\,\mathrm{Ortho}(X).$
Lemma. If $U\in O(n)$ is any orthogonal matrix, then $UE^{\prime\text{T}}$ has the same distribution as $E^{\prime\text{T}}$.
Proof. The rows of $E^{\prime}$ are i.i.d. Gaussian random vectors $\sim\mathcal{N}(0,\sigma^{2}I)$, which are isotropic. For each column, we have $E^{\prime}_{j}U^{\text{T}}\overset{d}{=}E^{\prime}_{j}$ and i.i.d. for different $j=1,\cdots,m$. Therefore, $E^{\prime}U^{\text{T}}\overset{d}{=}E^{\prime}$ and $UE^{\prime\text{T}}\overset{d}{=}E^{\prime\text{T}}$.
We now combine the lemmas to prove the main result: Let $U\in O(n)$ be any orthogonal matrix. Consider the distribution of $\mathrm{Ortho}(E^{\prime\text{T}}E)$ under the transformation induced by $U$. Using first lemma:
$\mathrm{Ortho}(UE^{\prime\text{T}}E)=U\,\mathrm{Ortho}(E^{\prime\text{T}}E).$
From the second lemma, since $UE^{\prime\text{T}}$ has the same distribution as $E^{\prime\text{T}}$, it follows that $UE^{\prime\text{T}}E$ has the same distribution as $E^{\prime\text{T}}E$. Consequently, $\mathrm{Ortho}(UE^{\prime\text{T}}E)$ has the same distribution as $\mathrm{Ortho}(E^{\prime\text{T}}E)$.
This invariance under left multiplication by any orthogonal matrix $U$ implies that $\mathrm{Ortho}(E^{\prime\text{T}}E)$ is uniformly distributed on $O(n)$ with respect to the Haar measure. ∎
Appendix E Inner Transformations in the Attention Module
Our main algorithm covers a basic set of inner transformations within attention. However, a more complete implementation might take into account the following complex relationship that could be contained within $W_{Q},W_{K},W_{V},W_{O}$. We present the extended version here, while the algorithm listing that we use in practice (see Appendix G) excludes some features like doing separate head permutation analysis.
The inner transformations for layer $\ell\ (1\leq\ell\leq L)$ is summarized as
$W_{Q,\ell}=\mu_{\ell}\cdot P_{1,\ell}\otimes P_{2,\ell}\otimes S_{\ell},\quad$
$W_{K,\ell}=\mu^{-1}_{\ell}\cdot P_{1,\ell}\otimes S_{\ell},$
$W_{V,\ell}=\sum_{v=1}^{\mathrm{NumKVHeads}}(\mathbf{1}_{v,\sigma_{\ell}(v)}\otimes H_{\ell,v}),$
$W_{O,\ell}=\sum_{v=1}^{\mathrm{NumKVHeads}}(\mathbf{1}_{v,\sigma_{\ell}(v)}\otimes P_{2}\otimes H_{\ell,v}),$
$(1\leq\ell\leq L)$